OpenStudy (anonymous):

How do you find the derivative of these problems?: h(x)=sin(2x)cos(2x) f(x)=cot(x)/(sin(x)) y=4sec^2x

OpenStudy (anonymous):

Product rule and quotient rule.

OpenStudy (anonymous):

(oh, and chain rule too)

OpenStudy (anonymous):

Yeah, but I'm having trouble using them to solve these. So for the first problem, this is what I did: h(x)=sin(2x)cos(2x) h'(x)=cos(2x)(cos(2x))(2)+(-sin(2x))(sin(2x)(2) That's as far as I made it to.

OpenStudy (turingtest):

and that's right

OpenStudy (turingtest):

now simplify

OpenStudy (anonymous):

I don't know how to simplify this though...

OpenStudy (anonymous):

sin(u) * sin(u) = sin^2(u)

OpenStudy (anonymous):

For the second one, remember that cot(u) = cos(u)/sin(u).